y=3x-1
and
y=negative2x
The x and y are the same in both problems.

Answer by checkley71(8403)

Y=3X-1 NOW SUBSTITUTE (3X-1) FOR Y IN THE OTHER EQUATION THUS
3X-1=-2X
3X+2X=1
5X=1
X=1/5 ANSWER
Y=3*1/5-1
Y=3/5-1
Y=-2/5 ANSWER